OpenStudy (anonymous):

find a general solution for dy/dt=(ty)^2

OpenStudy (danjs):

dy = (ty)^2 dt = t^2*y^2 dt

OpenStudy (danjs):

\[\frac{ 1 }{ y^2 }dy = t^2 dt\] integrate those

OpenStudy (danjs):

are you at the very beginning of DE class?

OpenStudy (anonymous):

this helps! but yes I am in the first week

OpenStudy (danjs):

have you done anything else besides separable equations

OpenStudy (danjs):

usually i think, substitutions and exact equations are next

OpenStudy (danjs):

been a couple years since that class

OpenStudy (anonymous):

no we haven't done much except basically the introduction. And she only touched on the separable equations that's why it's confusing to me

OpenStudy (danjs):

right, basically those are just equations where you can put each variable on its own side, then integrate to find the general function

OpenStudy (danjs):

It gets more fun later on

OpenStudy (danjs):

\[\int\limits \frac{ 1 }{ y^2 }dy =\int\limits t^2 dt\]

OpenStudy (danjs):

you get a function y(t)

OpenStudy (danjs):

\[\frac{ -1 }{ y } + C _{1} = \frac{ t^3 }{ 3 }+ C _{2}\]

OpenStudy (danjs):

You really dont need to remember ALL the integrating techniques, it helps, but basic U -Sub and Parts will get you through the course.

OpenStudy (anonymous):

yeah it seems easy enough I just need to practice. what do you think about dy/dt = 2y+1

OpenStudy (danjs):

c*e^(2x) - 1/2

OpenStudy (danjs):

just used my calc, that is the answer

OpenStudy (anonymous):

how do you do it so fast! -___- and how did you do it at all lol?

OpenStudy (danjs):

Oh i just used my calculator, for that one, since all you have done so far are separable equations, i assume it will be that

OpenStudy (anonymous):

yeah that's the answer in the back of the book I just don't see how you got that

OpenStudy (danjs):

\[dy = (2y + 1) dt \] \[\frac{ 1 }{ 2y+1 }dy = 1dt\]

OpenStudy (danjs):

integrate both sides

OpenStudy (danjs):

\[\frac{ \ln(2y+1) }{ 2 } = t + C\]

OpenStudy (danjs):

solve for y(t)

OpenStudy (danjs):

Both integrations produce a constant, C , you can combine the 2 into a single constant. Since it is arbitrary. Call it whatever you want.

OpenStudy (danjs):

I just put a C on the right side. C =C1-C2 where C2 was the constant from integrating the left side, and C1 the constant from the right side integration.

OpenStudy (danjs):

recall with exponents: \[e^a *e^b = e^{a+b}\]

OpenStudy (danjs):

Add me as a fan, and tag me with @DanJS when you get later into the course. It would be fun to bust out the old notebook and homework from DE and give them a go again.

OpenStudy (anonymous):

I think it's the solving for y(t) that I'm having a problem with. But I am new to the site so I'm just seeing how it's all working out. But I will DEFINITELY add you lol thank you for the help!

OpenStudy (danjs):

yeah, most of the work , like 80 or 90% will be algebra, expecially when you get to higher order DE, probably after the first test

OpenStudy (danjs):

Review partial fraction decompositions

OpenStudy (anonymous):

ahhh yes, I've forgotten all about that. that Christmas break away from math did me no good

OpenStudy (danjs):

I took a year off school after finishing calc 3 and linear algebra, i know the feeling, jumping into DE after a year break. It isn't too bad though. There arent many of those crazy integrals to work out. Most are straight forward.

OpenStudy (danjs):

just a ton of algebra

OpenStudy (danjs):

I would Recommend MIT OCW Differential Equations. Arthur Mattuk and MIT OCW Calculus Revisited (Black and White from like 1950) there are 5 lectures on DE's that deal with variation of parameters, and undetermined coefficient, that will be like next months work. ... ill find a link for you .. 1 sec

OpenStudy (danjs):

The first link is DE and Linear algebra together, some of the lectures are very helpful. The black and white ones, PART II (5 or so lectures) are a gold mine. Those help out a BUNCH when you get to those topics Those things got me an A in the course, and i understood it all very well

OpenStudy (anonymous):

@DanJS help me pleaseee

OpenStudy (anonymous):

yeah I'm about to go bookmark those now and look at them. I hope they help me as much as they helped you

OpenStudy (danjs):

I love the style of that old black and white teachers teaching. Easy to understand. Those will help out like next month when you get to second order DE's

OpenStudy (danjs):

cool i added you as a fan, tag me whenever you want @DanJS in a question.

OpenStudy (ribhu):

Arranging your equation as (dy/y^2)=tdt now integrating both the sides we get -1/y= (t^2)/2 + c where c is a constant of integration (t^2)/2+1/y+c=0 the final solution